Solve the following linear using the Addition Method. Describe in words the steps you will take to do that.
3x + 4y = -6
5x + 6y = -21/2

Answer by stanbon(75887)

3x + 4y = -6
5x + 6y = -21/2
--------------------
Multiply thru the 1st equation by 5 ; Multiply thru the 2nd equation by -3.
------------------------------
15x + 20y = -30
-15x - 18y = 63/2
-----------------------
Add so you can solve for "y":
2y = 3/2
y = 3/4
--------------
Substitute in 3x + 4y = -6 to solve for "x":
3x + 4(3/4) = -6
3x + 3 = -6
3x = -9
x = -3
-------------------------------
Cheers,
Stan H.